Problem 20.15 A helium ion (Q +2e) whose mass is 6.6x10 27 kg is accelerated by a voltage of 2300 V Part A What is its speed? Express your answer to two significant figures and include the appropriate units. Value Units Submit My Answers Give Up Part B What will be its radius of curvature if it moves in a plane perpendicular to a uniform 0.390 T field? Express your answer to two significant figures and include the appropriate units. Value Units Submit My Answers Give Up Part C What is its period of revolution? Express your answer to two significant figures and include the appropriate units T Value Units Submit My Answers Give Up
Answer #1

A.

speed, v = sqrt[2qV/m] = sqrt[2*3.2x10-19*2300/6.6x10-27] = 4.72x105 m/s

B.

radius: r = sqrt[2mV/qB2] = sqrt[2*6.6x10-27*2300 /3.2x10-19*(0.39)2] = 2.497x10-2 m = 2.5x10-2 m

C.

Period of revolution:

T = 2*pi*m/qB = 2*pi*6.6x10-27 / 3.2x10-19 * 0.39 = 3.32x10-7 s